OpenStudy (anonymous):

Would anyone be so kind to help me with Algebra 1? :)<3

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Ask away!

OpenStudy (anonymous):

Which equation is a quadratic equation? y + 3 = 5x + 10 y + 4x2 = (4x − 2)(x + 3) y − 8x = (x2 + 1)(9x − 7) y = 2x + 7x(x + 1)

OpenStudy (anonymous):

For sure?? Or pretty sure?

OpenStudy (anonymous):

let me check again :)

OpenStudy (anonymous):

K :)

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/quadratic-equation.html check out this website. it has ur answer :)

OpenStudy (anonymous):

So.. it's B??

OpenStudy (skullpatrol):

Don’t post only answers - guide the asker to a solution.

OpenStudy (usarmy3947):

its common sence u dont have to work it out and also its a word problem

OpenStudy (anonymous):

I'm confused..

OpenStudy (usarmy3947):

look he says like if the problem says two apples were found and 2 more apples where found he says we need to put 2+2=4 and the question is a equation

OpenStudy (anonymous):

So it's B?

OpenStudy (usarmy3947):

i guess so

OpenStudy (anonymous):

What? lol

OpenStudy (usarmy3947):

idk im that good in any thing just know when people are wrong

OpenStudy (anonymous):

Does anyone know what the answer is?

OpenStudy (usarmy3947):

if u want i can ask my friends? they know

OpenStudy (anonymous):

yeah

OpenStudy (usarmy3947):

ok ill ask on yahoo answers

OpenStudy (usarmy3947):

ok got to wait a few secs

OpenStudy (anonymous):

y - 8x = (x^2 + 1)(9x - 7) y - 8x = 9x^2 - 7x^2 + 9x - 7 y - 8x = 2x^2 + 9x - 7 y = 2x^2 + 9x + 8x - 7 y = 2x^2 + 15x - 7 I think it is C Because you want it in Ax^2 + Bx + C = 0

OpenStudy (usarmy3947):

y= 2x + 7x(x + 1)

OpenStudy (usarmy3947):

d.

OpenStudy (anonymous):

Question 17 (Multiple Choice Worth 5 points) [9.01] What is the value of the coefficient of "a" when the quadratic equation y = (3x + 1)(3x − 2) is written in standard form? −2 −3 9 6 -------------------------------------------------------------------------------- Question 18 (Multiple Choice Worth 5 points) [9.06] Using graphing technology, approximate the solutions of 0.25x2 − 0.46x − 3.1 = 0. x ≈ −2.72 and x ≈ 4.56 x ≈ −3.14 and x ≈ 4.24 x ≈ −2.88 and x ≈ 4.33 x ≈ −2.61 and x ≈ 4.84

OpenStudy (usarmy3947):

ur in flvs?

OpenStudy (anonymous):

yeah

OpenStudy (usarmy3947):

cool what city?

OpenStudy (usarmy3947):

can u help me in this question?

OpenStudy (usarmy3947):

An example of a shared power is to license professional workers raise and support armed forces control foreign trade provide health care and welfare